OpenStudy (anonymous):

given the guadratic function, y=3x^2+6x-2, find the vertex using the vertex formula x= -b over 2a A. (-1,-5) B. (2,-3) C. (-2,-3) D. (-1,5)

OpenStudy (anonymous):

The vertex is where the derivative is zero. \[\frac{dy}{dx}=6x+6=0\] Find this x. Substitute into original equation to find y. Done!

OpenStudy (amorfide):

okay you want to complete the square if you label your quadratic in terms of A B and C 3=A 6=B -2=C first of all you want to take out a factor of 3 3(x^2+2x-2/3) now we complete the square into (x+b/2)^2-(b/2)^2 3((x+1)^2 -1-2/3) multiply it all by 3 3(x+1)^2 -3-2 3(x+1)^2 -5 your vertex should be 1, -5

OpenStudy (amorfide):

that should be -1,-5